Here is the income statement for Ginsberg, Inc. GINSBERG, INC Income Statement For the Year Ended December 31, 2017 Sales revenue Cost of goods sold Gross profit Expenses (including $12,000 interest and $22,000 income taxes) Net income $400,000 250,000 150,000 100,000 $50,000 Additional information: 1. Common stock outstanding January 1, 2017, was 30,000 shares, and 40,000 shares were outstanding at December 31, 2017 2. The market price of Gillman, Inc., stock was $15.86 in 2017. 3. Cash dividends of $16,000 were paid, $4,500 of which were to preferred stockholders Compute the following measures for 2017. (Round to 2 decimal places, eg, 17.54% or 6.25: 1 or $2.66.) a) Earnings per share (b) Price-earnings ratio (c) Payout ratio (d) Times interest earned times times
Answer #1

Weighted average shares = 30000+(10000/2) = 35000 shares

a) Earning per share = Net income/Outstanding shares = 50000/35000 = 1.43 per share

b) Price earning ratio = MPS/EPS = 15.86/1.43 = 11.10 Times

c) Payout ratio = (16000-4500)/(50000-4500) = 25.27%

d) Time interest earned = (50000+12000+22000)/12000 = 7 Times